OpenStudy (anonymous):

PLEASSEEE HELP !!!!! Solve the equation.(work it out?) u + 2 = 4

OpenStudy (aroub):

u=2

OpenStudy (anonymous):

u + 2 = 4 -2 -2 u = 2